At a distance of 55 million light years from Earth, and with a mass equal to 2.4e12 solar masses, calculate the angular diameter of that black hole's event horizon as viewed from the Earth. (express your answer in arcseconds)

1
Expert's answer
2020-12-10T11:07:23-0500

The diameter of that black hole's event horizon (doubled Schwarzschild radius, see https://en.wikipedia.org/wiki/Schwarzschild_radius) is given as


"d_s = \dfrac{4GM}{c^2}"

where "G = 6.67\u00d710^{\u221211} m^3\u22c5kg^{\u22121}\u22c5s^{\u22122}" is the gravitational constant, "c = 3\cdot 10^8m/s" is the speed of light, and "M" is the mass of the black hole. Since it is equal to 2.4e12 solar masses, obtain:


"M = 2.4\times 10^{12}\cdot 1.99\times 10^{30}kg = 4.78\times 10^{42}kg"

Thus, obtain:

"d_s = \dfrac{4\cdot6.67×10^{−11}\cdot 4.78\times 10^{42}}{(3\cdot 10^8)^2} \approx 2.31\times 10^{16}m"

The angular diameter (in radians) is:


"\theta = \dfrac{d_s}{R}"

where "R" is the distance from the Earth to the black hole. Since this distance is 55 million light years, and one light year is approximately "9.46\times 10^{15}m", the distance is:


"R = 55\cdot 10^6\cdot 9.49\times 10^{15} \approx 5.2\times 10^{23}m"

Thus, the angular diameter is:


"\theta = \dfrac{2.31\times 10^{16}m}{5.2\times 10^{23}m} \approx 4.4423\times 10^{-8} rad"

To convert it to arcseconds, calculate:

"\theta =4.4423\times 10^{-8}\cdot 3600\cdot 180/\pi \approx 0.009''"

Answer. 0.009''.